OpenStudy (anonymous):

2x-4/x-5=0 How would I find x?

OpenStudy (anonymous):

I assume you mean 2x - (4/x) - 5 = 0. In that situation you would multiple both sides by x and then solve the polynomial equation.

OpenStudy (diyadiya):

\[\frac{ 2x-4 }{x-5 }=0 \ \ \ or \ \ 2x- \frac{4}{x}-5=0 \ \ \ or\ \ \ \frac{2x-4}{x}-5=0 \]

OpenStudy (diyadiya):

Which one is your question?

OpenStudy (goformit100):

Just rationalise it.